OpenStudy (anonymous):

Use calculus to locate the stationary points on the graph y=x+2/x

OpenStudy (anonymous):

I have located the x values of sqrt(2) and -sqrt(2) but i cannot solve that back into y=x+2x

OpenStudy (anonymous):

you got the right, what is the question?

OpenStudy (anonymous):

I cannot solve the x values into the original equation to find the y coordinates

OpenStudy (anonymous):

For the first one i have f'(sqrt(2))=Sqrt(2) + 2/sqrt(2)

OpenStudy (anonymous):

oh ok we can do that

OpenStudy (anonymous):

\[f(\sqrt2)=\sqrt2+\frac{2}{\sqrt{2}}=\sqrt{2}+\sqrt{2}=2\sqrt{2}\]

OpenStudy (anonymous):

How did you just get rid of the numerator 2 so easily

OpenStudy (anonymous):

same for the other one, except it is negative

OpenStudy (anonymous):

it is always the case that \(\frac{a}{\sqrt{a}}=\sqrt{a}\)

OpenStudy (anonymous):

How is that? I have never learnt that before?

OpenStudy (anonymous):

rationalize the denominator, and you will see it instantly

OpenStudy (anonymous):

but it is \[\sqrt{2} + 2/\sqrt{2}\]

OpenStudy (anonymous):

or if you prefer, show \[\left(\frac{a}{\sqrt{a}}\right)^2=a\]

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

and since \(\frac{2}{\sqrt{2}}=\sqrt{2}\) you get \[\sqrt{2}+\sqrt{2}=2\sqrt{2}\]

OpenStudy (anonymous):

How is sqrt(2) + sqrt(2) = 2*sqrt(2)

OpenStudy (anonymous):

oh bugger its like terms. sorry im not thinking

OpenStudy (anonymous):

Sorry, thanks for the help mate, i think i get it now

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Have a good one :)